OpenStudy (anonymous):

Arc AB of circle O contains 72 degrees and is 6 pi long. Find area of circle O.

OpenStudy (paxpolaris):

first find the radius from the length of the Arc

OpenStudy (anonymous):

What's the relationship between an arc and radius?

OpenStudy (anonymous):

@PaxPolaris ?

OpenStudy (paxpolaris):

length of arc \[=\frac \theta{360}\times 2\pi r\]

OpenStudy (paxpolaris):

if \(\theta\) is in degrees

OpenStudy (anonymous):

r= 15

OpenStudy (anonymous):

Is the answer 225 pi?

OpenStudy (paxpolaris):

yes :)

OpenStudy (anonymous):

tysm~